a positively charged particle p enters the region between two parallel plates with a velocity u in the direction parallel to the plate there is a uniform electric field in this region p emergers from this region with a velocity v take c as a constant v will depend on u as 

Grade:12th pass

1 Answers

Arun
3 years ago
Let l be the length of each plate and ux = u = constant. ∴ the time of travel between the plates is given by t = l/u. Let a be the constant acceleration in the y-direction. ∴ vy = at when the particle emerges from the plates. ∴ v2 = ux2 + vy2 = u2 + a2t2 = u2 + a2 l2/u2 = u2 + α/u2.